OpenStudy (ksaimouli):

let p(1,3,2) and let L be the line with parametric equation x=2-t,y=-1+2t, z=3+t use the cross product to find the distance between P to L

OpenStudy (ksaimouli):

I have no idea, but I tried to find the plane that passes thru the line and used point and plane distance formula.

OpenStudy (science0229):

Look at your parametric equation. Can you put it into a form\[\frac{ x-x _{1} }{ a }=\frac{ y-y _{1} }{ b }=\frac{ z-z _{1} }{ c }=t\]

OpenStudy (science0229):

hello?

OpenStudy (ksaimouli):

yup, \[\frac{ x-2 }{ -1 }=\frac{ y+1 }{ 2 }=\frac{ z-3 }{ }\]

OpenStudy (ksaimouli):

1

OpenStudy (ksaimouli):

@science0229

OpenStudy (science0229):

right!

OpenStudy (science0229):

Now, in that form, we know that the vector form is v=<a,b,c>, or <-1,2,1>

OpenStudy (science0229):

Wait, do you know the symmetrical line equation?

OpenStudy (ksaimouli):

ys

OpenStudy (science0229):

ok. then does this make sense to you so far?

OpenStudy (ksaimouli):

yes

OpenStudy (science0229):

Now, the formula for the distance between plane 'n' and line 'v' is known as \[\frac{ n*v }{ |n| }\] * should be a dot.

OpenStudy (ksaimouli):

is n= (1,3,2)

OpenStudy (science0229):

yes

OpenStudy (ksaimouli):

\[\frac{ 7 }{ \sqrt{14}}\]

OpenStudy (ksaimouli):

"Now, the formula for the distance between plane 'n' and line 'v' is known as" n is not the plane it is a point

OpenStudy (science0229):

the answer is correct

OpenStudy (ksaimouli):

no, \[\it is \frac{ \sqrt{66} }{ 3}\]

OpenStudy (ksaimouli):

it is that

OpenStudy (science0229):

Wait. Oops. I thought p meant plane :(

OpenStudy (ksaimouli):

P is the point

OpenStudy (phi):

I think the idea is this